OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 320, 80, 20, 5,...

OpenStudy (anonymous):

Give the answer

OpenStudy (anonymous):

320 time 1/4 gives 80 times 1/4 gives 20 ratio=1/4 intial value = a summation: sigma a / 1-r = 320/.75

OpenStudy (anonymous):

or if you want terms, 1/4b = c 1/4c = d etc..

OpenStudy (anonymous):

hope that helps!